A chemist wishes to mix some pour acid with some water to peoduce 16 L of a solution that is 30% acid. How much pure acid and how much water should be mixed?
Answer by josmiceli(19441)

a = amount of acid
w = amount of water
[1] a + w = 16 L (given)
[2] a / w = .3
from [2]
a = .3w
substitute in [1]
.3w + w =16
1.3w = 16
w = 160/13 = 12.31
a = (3/10)(160/13) = 48/13 = 3.69
check
12.31 + 3.69 = 16
3.69 / 12.31 = .3